A friend devises a game that is played by rolling a single six-sided die once. If you roll a 6, he pays you $3; if you roll a 5, he pays you nothing; if you roll a number less than 5, you pay him $1.
Answer #1

From the given data, the following Table is calculated:

Event Value (x) Probability (p) x p
Roll a 6 3 1/6 3 X 1/6 = 0.50
Roll a 5 0 1/6 0X 1/6 = 0.00
Roll 1,2,3,4 - 1 4/6 - 1 X 4/6 = - 0.67
Total = - 0.17

So,

Expected Value of the game = - 0.17

Since the expected value of the game is negative, the conclusion is that playing the game is loss for you and it is profitable for your friend.